OpenStudy (anonymous):

How many solutions does the equation 5x + 3x − 4 = 10 have? Zero One Two Infinitely many

OpenStudy (anonymous):

@aryandecoolest @AaronAndyson @autumnm @Adi3 @HELPMEPLZ!! @HelpOfTheGods @CallMeKiki @courtneygraley009 @DarkMoonZ @Loser66

OpenStudy (welshfella):

what do you think?

OpenStudy (anonymous):

@michayla223 @GreatGuy @minisweet4 @baconbaby1334 @selfconfidence777

OpenStudy (anonymous):

i think it is b

OpenStudy (welshfella):

try solving it

OpenStudy (anonymous):

i have i got no were i could not find were it = 10 but i now its not a

OpenStudy (welshfella):

why?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

its zero

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

i tryed that and got it wrong

OpenStudy (welshfella):

5x + 3x − 4 = 10 add 4 to both sides 5x + 3x = 10 + 4 add like terms 8x = 14 so x = ?

OpenStudy (anonymous):

how is it not a